OpenStudy (anonymous):

create a cubic polynomial inequality for which x=3 or x<=-2 is the solution. explain your reasoning

OpenStudy (anonymous):

what does that last one mean?

OpenStudy (anonymous):

oh it is an inequality so what does the first one mean?

OpenStudy (anonymous):

\[\le\]

OpenStudy (anonymous):

so it is \[x\leq -2\] or \[x\leq 3\]?? this doesn't really make much sense since takent together they mean \[x\leq 3\]

OpenStudy (anonymous):

x=3 or x\[\le\]2

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

sorry i don't get it. one is an inequality, one is an equality

OpenStudy (anonymous):

I don't understand the question :(

OpenStudy (anonymous):

can you write down exactly what it says?